# SI_appendix_phys

### v1

@@ Abstract
The formation of dynamical patterns is one of the most striking features of non-equilibrium physical systems. Recent work has shown that such patterns arise generically from forces that violate Newton’s third law, known as non-reciprocal interactions. These non-equilibrium phenomena are challenging for modern theories. Here, we introduce a model mixture of active (self-propelled) and passive (diffusive) particles with non-reciprocal effective interactions, which is amenable to exact mathematical analysis. We exploit state-of-the-art methods to derive exact hydrodynamic equations for the particle densities. We study the resulting collective behavior, including the linear stability of homogeneous states and phase coexistence in large systems. This reveals a novel phase diagram with the spinodal associated with active phase separation protruding through the associated binodal, heralding the emergence of dynamical steady states. We analyze these states in the thermodynamic limit of large system size, showing, for example, that sharp interfaces may travel at finite velocities, but traveling phase-separated states are forbidden. The model’s mathematical tractability enables precise new conclusions beyond those available by numerical simulation of particle models or field theories.
@@ Introduction
Simple systems of interacting particles (or agents) can support complex emergent behavior, including the self-assembly of nanoscale equilibrium structures, the self-organization of animals into flocks and swarms, and pattern formation in chemical reactions. Describing these effects has been a long-standing challenge for physics and mathematics: modern theories focus on emergent non-equilibrium behavior, which continues to challenge our understanding, and our ability to predict macroscopic collective phenomena from the underlying microscopic rules. Recent studies have highlighted that non-reciprocal interactions in non-equilibrium systems lead generically to pattern formation in a variety of physical settings [1], including reaction-diffusion systems [2, 3, 4], living chiral crystals [5, 6] and quorum sensing bacteria [7, 8]. Despite their diversity, these systems appear to self-organize according to a common set of physical principles, offering the opportunity for a predictive theory with broad scope.
@@
Within this context, the non-reciprocal Cahn–Hilliard (NRCH) equation has recently emerged as a canonical model for non-reciprocally coupled particle models [9, 10, 11]. It illustrates that phase-separated systems can undergo exceptional phase transitions when subjected to non-reciprocal driving, leading to dynamical (traveling) steady states. This simple and elegant equation bridges the established equilibrium theory of phase separation and the complex world of pattern formation in non-equilibrium systems. However, pattern formation in non-reciprocally coupled systems continues to challenge our understanding, including the vital question of which patterns will appear in any given system [12, 13, 1, 14].
@@
This work addresses these challenges by analyzing a specific non-reciprocal system for which exact mathematical results can be derived. Specifically, we introduce an idealized mixture of interacting active and passive particles and derive its hydrodynamic limit, the equation that governs its large-scale collective behavior. Such mixtures are known to display many features of non-reciprocally interacting systems and NRCH-like equations that approximate the motion of these mixtures have been proposed, either on phenomenological grounds or by various approximation schemes [15, 16, 9]. These capture many qualitative features of the resulting dynamics but are only partially quantitative. Our exact hydrodynamic equation differs from NRCH, but the resulting phenomenology is similar and consistent with generic principles of self-organization via non-reciprocity.
In the mixture, active particles alone result in stationary, motility-induced phase separation (commonly referred to as MIPS [17]), but adding an extra population of passive (diffusing) particles is sufficient to induce patterns, including traveling clusters, where self-organized groups of active particles push their passive counterparts around the system [18, 19]. By studying the hydrodynamic equation in large domains, we relate traveling interfaces to those of static MIPS clusters, providing a new link between the equilibrium-like physics of phase separation and the dynamical patterns characteristic of non-reciprocity.
@@
All these results rely on state-of-the-art mathematical techniques together with the idealized underlying model. In particular, we assume that particles move on an underlying two-dimensional lattice with stochastic dynamical rules and that self-propulsion only occurs in the left and right directions. These idealized features enable a rigorous hydrodynamic limit [20, 21]: when the lattice spacing tends to zero and the number of particles to infinity, the particle densities obey deterministic continuum equations, which we derive exactly. The resulting system differs from NRCH: it has some similar features but also reveals interesting new behavior. These effects are especially pronounced when we analyze large systems, such that the system can phase-separate into dense liquid and dilute vapor regions, separated by narrow interfaces. For these large systems, we characterize phase-separated and pattern-forming states by an asymptotic method that separates narrow interfacial regions from the bulk of the system. Our results include an unusual type of phase diagram where the spinodal curve for the liquid-vapor transition protrudes through the binodal, signaling the onset of pattern formation. We also demonstrate the existence of asymmetric traveling patterns that feature sharp interfaces, and patterns formed of counterpropagating clusters.
@@
Our microscopic model clearly exhibits the principles of non-reciprocal self-organization. Moreover, its idealized features enable us to draw sharp conclusions about large length and time scales, which would be extremely challenging to obtain from numerical simulations of more complicated models. Our analysis of the large-system limit shows how ideas coming from equilibrium phases and their interfaces can be carried over to pattern-forming states; they also enable a fully non-linear treatment of the pattern-forming (traveling) steady states, illustrating much more complex behavior than could be predicted by linear stability analysis of the homogeneous state. These exact results mean that our idealized model serves as a reference point against which future non-reciprocal systems can be compared.
I Active-Passive Lattice Gas (APLG) Model
@@
Figure 1: Phase diagrams spanned by the active and passive particles’ volume fractions $\phi_{a}$ and $\phi_{p}$, respectively, for (a) $\text{Pe}=5$ and (b) $\text{Pe}=7.5$. The spinodal (blue curve) encloses the region of linear stability of homogeneous solutions. The binodal (red curve) encloses the region of phase separation. Intersection points of the tie lines (gray lines) with the binodal indicate the composition of the liquid and vapor phases. Black crosses indicate bifurcations of co-dimension two; see text for discussion. Regions with stable Homogeneous (H), Phase Separated (PS), and Dynamic (D) steady states are marked on the diagram. [The white region in (b) is left unlabelled because the boundary between H+D and H is unknown.]
@@
It is convenient to rescale time and space by $D_{R}^{-1}$ and $\sqrt{D_{T}/D_{R}}$ respectively, and introduce the Péclet number $\text{Pe}=v_{0}/\sqrt{D_{T}D_{R}}$. A configuration of the APLG is defined in terms of occupancies: $\eta_{\sigma}(\bm{x},t)\in\{0,1\}$ is the number of particles of type $\sigma$ at site $\bm{x}$ and time $t$. The hydrodynamic limit equations describe the evolution of the local densities $\rho_{\sigma}(\bm{x},t)$ of particles of type $\sigma\in\{+1,-1,0\}$, as the lattice spacing $h\to 0$. Building on [21], we rigorously derive the hydrodynamic limit of the APLG model, obtaining exact macroscopic evolution equations for the densities $\rho_{\sigma}$,
@@
${\cal D}(\rho)=[1-d_{s}(\rho)]/\rho,\quad s(\rho)={\cal D}(\rho)-1,$ (2)
see Appendix A for details. Technically, the APLG model is of non-gradient type in the sense of [20], which means that the mean-field approximation $\langle\eta_{\sigma}(\bm{x},t)\eta_{\sigma}(\hat{\bm{x}},t)\rangle\approx\rho_{\sigma}(\bm{x},t)\rho_{\sigma}(\hat{\bm{x}},t)$ does not hold and no explicit formula for $d_{s}(\rho)$ exists. However, it is very accurately approximated by [25]:
@@
where $\alpha=\pi/2-1$ in two dimensions [for other accurate approximations, see [26, 27, 28, 29, 30, 23, 25, 31]]. Here, we use (3) in the numerical analysis of the exact hydrodynamic limit (1).
Mathematical analysis of (1) is challenging due to the density dependence of the coefficients. However, since the active self-propulsion is only in the horizontal ($x$) direction, the equations are diffusive in the vertical direction [24], so that any variation with respect to $y$ will converge to zero over time, and instabilities of the homogeneous state are also independent of $y$. We exploit this symmetry throughout, restricting solutions of (1) to the form $\rho_{\sigma}(x,y,t)=\rho_{\sigma}(x,t)$. Four dimensionless parameters govern these solutions: the Péclet number Pe, the rescaled domain length in the horizontal axis $L=\ell_{x}\sqrt{D_{T}/D_{R}}$, the active volume fraction $\phi_{a}$, and passive volume fraction $\phi_{p}$.
II Overview of phase behavior
The APLG model supports different dynamical phases. We focus on behavior in the thermodynamic limit of large domains ($L\gg 1$), which allows some properties to be derived analytically. As in pure active systems, [23], the APLG supports phase-separated (PS) stationary states with macroscopic dense (liquid) and dilute (vapor) regions. The densities of active and passive particles in the liquid and vapor phases trace out the binodal curve in the ($\phi_{a},\phi_{p})$ plane, see Fig. 1(a). For $L\gg 1$, this curve can be calculated numerically exactly following the prescription of [32]. This relies on the observation that, for PS states, local concentrations of passive particles and vacancies are proportional
@@ II Overview of phase behavior
We also compute the spinodal curve, as the limit of stability of the homogeneous state (which is $\rho_{\pm}=\phi_{a}/2$, $\rho_{0}=\phi_{p}$). Linear stability analysis of (1) shows that the homogeneous state is unstable when Pe is sufficiently large and $\phi_{p}$ is sufficiently small (see Methods). For the case shown in Fig. 1(a), the dominant eigenvalue of the stability problem is always real, as in the pure active case. Then, the binodal and spinodal are tangent at the critical point, which is the standard phenomenology of liquid-vapor phase coexistence for two-component mixtures, see Fig. 1(a). [Inside the spinodal, the homogeneous state is linearly unstable, and the phase-separated (PS) state is stable; between the spinodal and binodal, the PS state is globally stable while the homogeneous state is metastable.]
For larger Pe, this familiar scenario changes qualitatively. In particular, (1) includes non-reciprocal couplings between active and passive densities. As a result, the dominant eigenvalues of the linear stability problem may become complex [9, 19, 15], which typically leads to dynamical steady states. Such a scenario is illustrated in Fig. 1(b). Crosses on the spinodal mark co-dimension two (Bogdanov–Takens) bifurcations [33] where the dominant eigenvalue becomes complex and the instability of the homogeneous state changes from a pitchfork to a Hopf bifurcation [34]. Importantly, the binodal curve can still be computed for this system: we observe that the spinodal protrudes through the binodal. This effect is intrinsically linked to the existence of complex eigenvalues (Appendix, section C) and signals the onset of new physics.
Figure 2: (a),(b) Kymographs showing the spatiotemporal dynamics of a phase-separating solution. (a) $y$-averaged local density of a particle simulation (see Methods). (b) Numerical solution to (1), the initial condition is a homogenous state with a uniform random perturbation (see Appendix 1 for details). (c) Density profile at $t=20$. Black lines display the solution to (1), and red lines display the local density of the particle simulation. Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.5$, $\phi_{p}=0.1$, $\Delta x=0.01$ (PDE model), $h=0.01$ (particle model).
Within the protruding part of the spinodal, the homogeneous state is linearly unstable, and PS states do not exist. It follows that some dynamical steady states must be present in this region. Moreover, in regions where PS states do exist, it may be that (at least) one of the coexisting phases is linearly unstable, which also renders the PS state unstable. The result is that dynamical steady states must exist throughout the blue-shaded region D in Fig. 1(b); these will be investigated in detail below. Note that this does not rule out the existence of dynamical states elsewhere. Also, while the dominant eigenvalue for the instability is complex for a large part of the spinodal in Fig. 1(b), the resulting steady state may still be (stationary) PS, showing that steady-state properties cannot be deduced directly from linear stability analysis.
@@
Figure 3: (a),(b) Kymographs showing the spatiotemporal dynamics of a CP solution. (a) The $y$-averaged local density of a particle simulation (see Methods). (b) The solution to (1) whose initial condition is a homogenous state with a uniform random perturbation (see Appendix 1). (c) Density profiles before, during, and after the collision of two counter-propagating interfaces ($t=17,18,19,20$). Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.3$, $\phi_{p}=0.6$, $\Delta x=0.01$, $h=0.01$.
III Illustrative steady states
We present numerical results that illustrate the behavior of the APLG, including direct simulation of the particle model by the Gillespie algorithm [36], and time-stepping the (deterministic) partial differential equation (1). We consider fairly small domains so that the particle model simulations are tractable. (The total number of particles is approximately $(\ell_{y}/\ell_{x})\phi(L/h)^{2}$ and the simulations of Figs. 2 and 3 already involve thousands of particles.)
@@ III Illustrative steady states
The binodal construction (Fig. 1) demonstrates the existence of stationary solutions to (1) for large system size $L$. These consist of large $O(L)$ liquid and vapor regions, separated by interfaces of $O(1)$ width. In the pure active case ($\phi_{p}=0$), this is motility-induced phase separation (MIPS) [17, 37, 38, 39], which is also a familiar phenomenon in non-reciprocal matter [9, 10].
Fig. 2 shows results of particle-based simulations of the APLG and corresponding numerical solutions of (1) in a domain of size $L=2$. Phase separation occurs in both cases, starting from homogeneous initial states. For these parameters, the dense phase is dominated by active particles, while the dilute phase is mostly passive. In fact, denser phases always have lower concentrations of passive particles because of (4). As usual for MIPS, the magnetization $m$ is large in the interfacial regions but very small within the phases. Note that while the analytic binodal computation shows that stationary PS states exist, these numerical results also show that they are stable (for these parameters).
Figure 4: Traveling profile solution to (1), whose initial condition is the homogeneous state perturbed by left and right traveling unstable sinusoidal modes and a uniform random perturbation (see Appendix 1 for details). (a) Density profile at $t=500$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=25.0$, $\phi_{a}=0.36$, $\phi_{p}=0.3$, $\Delta x=0.05$.
@@
We now turn to dynamical steady states, which appear in the (blue) region D in Fig. 1(b). We focus on two types of state: CP solutions retain an overall left-right symmetry with clusters of particles moving in both directions; T solutions break left-right symmetry, leading to a density profile that travels at a fixed speed, specifically
$\rho_{\sigma}(x,t)=\varrho_{\sigma}(x-ct/L)$ (6)
@@
where $c$ is a constant so the wave velocity is $c/L$, the reason for this $L$-dependence will be discussed below. (There is an analogy of CP and T solutions with standing waves and traveling waves respectively, but note that both CP and T solutions are strongly anharmonic.)
Figure 5: T solutions for the large-$L$ problem (8) and the finite-$L$ problem (7) at $\text{Pe}=7.5$. (Top) Solutions to (8) as a function of $z/L$. (Bottom) Solutions plotted parametrically on the phase diagram of Fig. 1(b), with solid lines for solutions of (8) and dashed lines for solutions of (7). Black crosses mark the volume fractions (spatially-averaged densities). Densities: $(\phi_{a},\phi_{p})=(0.45,0.35),(0.30,0.25),(0.55,0.20),(0.538,0.20)$ for (a,b,c,d) respectively. Other parameters: $N=1024$ for numerical solution of (8) and $L=25$, $\Delta z=0.015625$ for (7). Wave speeds for (a,b,c,d) are $c=2.0091,2.8979,0.3526,0.2198$ (4 decimal places).
@@
An example CP state is shown in Fig. 3, which again compares direct simulation of the particle model with the numerical solution of (1) whose initial condition is a homogenous state with a uniform random perturbation (see Appendix 1 for details). This perturbation grows via an instability that involves counterpropagating sinusoidal waves with equal speeds and growth rates. After this transient growth, the resulting time-periodic state consists of two oppositely polarised active clusters that move in opposite directions. As they move, they accumulate passive particles in front of them, via a “snowplow effect”. The clusters slow down during collisions, but they eventually pass through each other, and the cycle continues. Similar time-periodic states have been observed in other active and non-reciprocal systems [40, 11, 41, 42]: their presence here emphasises that they are generic. Despite the analogy with standing waves, we emphasize that these CP clusters experience complex (nonlinear) scattering processes when they meet (see for example Fig. S3).
@@
An illustrative T state is shown in Fig. 4, obtained by time-stepping (1) for a larger system ($L=25$). After an initial transient, we find a solution of the form of (6). As in the CP case, this consists of a localized packet of active particles that pushes passive particles in front of it. We find numerically that such solutions can have a variety of shapes and non-trivial dependence on system size. To simplify this diverse behavior, we again turn to large systems ($L\gg 1$), which enables analytic progress, by analogy with PS states. Hence we reveal new connections between dynamical patterns in this system and equilibrium phase coexistence.
IV Traveling solutions in large systems

### v2

# Dynamical patterns and nonreciprocal effective interactions in an active-passive mixture through exact hydrodynamic analysis

@@ Abstract
The formation of dynamical patterns is one of the most striking features of nonequilibrium physical systems. Recent work has shown that such patterns arise generically from forces that violate Newton’s third law, known as nonreciprocal interactions. These nonequilibrium phenomena are challenging for modern theories. Here, we introduce a model mixture of active (self-propelled) and passive (diffusive) particles amenable to exact mathematical analysis. We exploit state-of-the-art methods to derive exact hydrodynamic equations for the particle densities, which reveal effective nonreciprocal couplings between the active and passive species. We study the resulting collective behavior, including the linear stability of homogeneous states and phase coexistence in large systems. This reveals a novel phase diagram with the spinodal associated with active phase separation protruding through the associated binodal, heralding the emergence of dynamical steady states. We analyze these states in the thermodynamic limit of large system size, showing, for example, that sharp interfaces may travel at finite velocities, but traveling phase-separated states are forbidden. The model’s mathematical tractability enables precise new conclusions beyond those available by numerical simulation of particle models or field theories.
11footnotetext: bruna@maths.ox.ac.uk22footnotetext: These authors contributed equally to this work.
@@ Introduction
Simple systems of interacting particles (or agents) can support complex emergent behavior, including the self-assembly of nanoscale equilibrium structures, the self-organization of animals into flocks and swarms, and pattern formation in chemical reactions. Describing these effects has been a long-standing challenge for physics and mathematics: modern theories focus on emergent nonequilibrium behavior, making it difficult to predict macroscopic collective phenomena from the underlying microscopic rules. Recent studies have highlighted that nonreciprocal interactions in nonequilibrium systems lead generically to pattern formation in a variety of physical settings [1], including reaction-diffusion systems [2, 3, 4], living chiral crystals [5, 6] and quorum sensing bacteria [7, 8]. Despite their diversity, these systems appear to self-organize according to a common set of physical principles, offering the opportunity for a predictive theory with broad scope.
@@
Within this context, the nonreciprocal Cahn–Hilliard (NRCH) equation has recently emerged as a canonical model for nonreciprocally coupled particle models [9, 10, 11]. It illustrates that phase-separated systems can undergo exceptional phase transitions when subjected to nonreciprocal driving, leading to dynamical (traveling) steady states. This simple and elegant equation bridges the established equilibrium theory of phase separation and the complex world of pattern formation in nonequilibrium systems. However, pattern formation in nonreciprocally coupled systems continues to challenge our understanding, including the vital question of which patterns will appear in any given system [12, 13, 1, 14].
@@
This work addresses these challenges by analyzing a specific nonreciprocal system for which exact mathematical results can be derived. Specifically, we introduce an idealized mixture of interacting active and passive particles and derive its exact hydrodynamic limit, the equation that governs its large-scale collective behavior. Such mixtures are known to display many features of nonreciprocally interacting systems and NRCH-like equations that approximate their motion have been proposed, either on phenomenological grounds or by various approximation schemes [15, 16, 9, 17]. These capture many qualitative features of the resulting dynamics but are only partially quantitative. Our hydrodynamic equation leads to phenomenology similar to that of NRCH and is consistent with generic principles of self-organization via nonreciprocity.
@@
Our analysis relies on state-of-the-art mathematical techniques together with a simple underlying model. In particular, we assume that particles move on an underlying two-dimensional lattice with stochastic dynamical rules and that self-propulsion only occurs in the left and right directions. These idealized features enable a rigorous hydrodynamic limit [18, 19]: when the lattice spacing tends to zero and the number of particles to infinity, the particle densities obey deterministic continuum equations, which we derive exactly. The resulting system differs from NRCH: it has some similar features but also reveals interesting new behavior which we outline next.
In our model, active particles alone result in stationary, motility-induced phase separation (commonly referred to as MIPS [20]), but adding an extra population of passive (diffusing) particles can induce patterns, including traveling clusters, where self-organized groups of active particles push their passive counterparts around the system [21, 22], see also [23]. The hydrodynamic equation reveals an unusual type of phase diagram where the spinodal curve for the liquid-vapor transition protrudes through the binodal, signaling the onset of pattern formation. In this region we demonstrate the existence of asymmetric traveling patterns and patterns formed of counterpropagating clusters. For large systems, these patterns feature sharp interfaces that separate dense liquid and dilute vapor regions. By studying the hydrodynamic equation in large domains, we relate traveling interfaces to those of static MIPS clusters, providing a new link between the equilibrium-like physics of phase separation and the dynamical patterns characteristic of nonreciprocity.
@@
Our study provides an explicit and important example of how nonreciprocal effective interactions can emerge at the macroscale from simple interactions at the microscale. Specifically, while the direct interactions among particles are reciprocal – resulting solely from volume exclusion – particle simulations and analyses of the hydrodynamic equation clearly demonstrate that the model exhibits the principles of nonreciprocal self-organization. Furthermore, its idealized features allow us to draw sharp conclusions about large length and time scales, which would be extremely challenging to achieve from numerical simulations of more complex models. Our analysis of the large-system limit demonstrates how concepts from equilibrium phases and their interfaces can be applied to pattern-forming states. These features also facilitate a fully nonlinear treatment of the pattern-forming (traveling) steady states, revealing much more intricate behavior than could be predicted by linear stability analysis of the homogeneous state.
Based on these exact results, we identify several phenomena that should be considered as reference points for future studies of nonreciprocal systems. Specifically: the protrusion of a spinodal line through the binodal is a generic mechanism for instability of static phase separation, leading to dynamical patterns; in large systems, such patterns may incorporate interfaces between dense and dilute regions resembling those found in static phase separation; the resulting patterns may travel with a fixed velocity or they may include different traveling objects that interact repeatedly. These results significantly expand the phenomenology of nonreciprocally coupled mixtures. The separation of length scales between sharp interfaces and macroscopic dense/dilute domains suggests focusing on dynamical patterns in a suitable large-system limit in order to draw useful analogies with static (thermodynamics) phase transitions. We discuss these points further in later sections.
Results
.1 Active-Passive Lattice Gas (APLG) Model
@@ Results
Figure 1: Phase diagrams spanned by the active and passive particles’ volume fractions $\phi_{a}$ and $\phi_{p}$, respectively, for (a) $\text{Pe}=5$ and (b) $\text{Pe}=7.5$. The spinodal (blue curve) encloses the region of linear stability of homogeneous solutions. The binodal (red curve) encloses the region of phase separation. Intersection points of the tie lines (gray lines) with the binodal indicate the composition of the liquid and vapor phases. Black crosses indicate bifurcations of co-dimension two; see text for discussion. Regions with stable Homogeneous (H), Phase Separated (PS), and Dynamic (D) steady states are marked on the diagram. [In (b), the white region is left unlabelled because the boundary between H+D and H is unknown and the grey region corresponds to $\phi_{a}+\phi_{p}>1$, which is inaccessible due to size exclusion. The black tie line intersects the spinodal, so phase-separated states above this line are linearly unstable.]
@@
The number of lattice sites in the APLG is $\ell_{x}\ell_{y}/h^{2}$, and we identify the lattice spacing $h$ with the size of a particle. The hydrodynamic limit is $h\to 0$ at fixed $\ell_{x},\ell_{y},\phi_{a},\phi_{p}$: it corresponds to “zooming out”, in order to describe motion on scales much larger than the particle size. The APLG dynamics above has nontrivial $h$-dependence [for example, the hopping rate is $O(h^{-2})$ but the orientation flip rate is $O(1)$]. This choice of scaling limit enables rigorous mathematical calculations [19]. Other types of large-system limit are discussed in Appendix 1.
To take the hydrodynamic limit, it is convenient to rescale time and space by $D_{R}^{-1}$ and $\sqrt{D_{T}/D_{R}}$ respectively, and introduce the Péclet number $\text{Pe}=v_{0}/\sqrt{D_{T}D_{R}}$. A configuration of the APLG is defined in terms of occupancies: $\eta_{\sigma}(\bm{x},t)\in\{0,1\}$ is the number of particles of type $\sigma$ at site $\bm{x}$ and time $t$. The hydrodynamic limit equations describe the evolution of the local densities $\rho_{\sigma}(\bm{x},t)$ of particles of type $\sigma\in\{+1,-1,0\}$, as the lattice spacing $h\to 0$. Building on [19], we rigorously derive the hydrodynamic limit of the APLG model, obtaining exact macroscopic evolution equations for the densities $\rho_{\sigma}$ (see Appendix A)
@@
${\cal D}(\rho)=[1-d_{s}(\rho)]/\rho,\quad s(\rho)={\cal D}(\rho)-1.$ (2)
[…]
@@
with $\alpha=\pi/2-1$, which exactly matches the value and the first derivative of $d_{s}(\rho)$ at $\rho=0,1$, and approximates $d_{s}$ very accurately for all $\rho$. For other accurate approximations, see [30, 31, 32, 33, 34, 35]. Here, we use (3) in the numerical analysis of the exact hydrodynamic limit (1). Corrections to the mean-field approximation $\langle\eta_{\sigma}(\bm{x},t)\eta_{\sigma}(\hat{\bm{x}},t)\rangle\approx\rho_{\sigma}(\bm{x},t)\rho_{\sigma}(\hat{\bm{x}},t)$ are important in the APLG model and enter the hydrodynamic limit; the mean-field approximation would lead to $d_{s}(\rho)=(1-\rho)$. The method to obtain (1) is valid for any dimension greater than one.
Mathematical analysis of (1) is challenging due to the density dependence of the coefficients. However, since the active self-propulsion is only in the horizontal ($x$) direction, the equations are diffusive in the vertical direction [26], so that any variation with respect to $y$ will converge to zero over time, and instabilities of the homogeneous state are also independent of $y$. We exploit this symmetry throughout, restricting solutions of (1) to the form $\rho_{\sigma}(x,y,t)=\rho_{\sigma}(x,t)$. Four dimensionless parameters govern these solutions: the Péclet number Pe, the rescaled domain length in the horizontal axis $L=\ell_{x}\sqrt{D_{R}/D_{T}}$, the active volume fraction $\phi_{a}$, and passive volume fraction $\phi_{p}$.
.2 Overview of phase behavior
Figure 2: (a),(b) Kymographs showing the spatiotemporal dynamics of a phase-separating solution. (a) $y$-averaged local density of a particle simulation (see Methods). (b) Numerical solution to (1), the initial condition is a homogenous state with a uniform random perturbation (see Appendix 1 for details, Eq. (S62)). (c) Density profile at $t=20$. Black lines display the solution to (1), and red lines display the local density of the particle simulation. Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.5$, $\phi_{p}=0.1$, $\Delta x=0.01$ (PDE model), $h=0.01$ (particle model).
Figure 3: (a),(b) Kymographs showing the spatiotemporal dynamics of a CP solution. (a) The $y$-averaged local density of a particle simulation (see Methods). (b) The solution to (1) whose initial condition is a homogenous state with a uniform random perturbation (see Appendix 1, Eq. (S62)). (c) Density profiles before, during, and after the collision of two counter-propagating interfaces ($t=17,18,19,20$). Parameters: $\text{Pe}=20$, $L=2$, $\phi_{a}=0.3$, $\phi_{p}=0.6$, $\Delta x=0.01$, $h=0.01$.
The APLG model supports different dynamical phases. We focus on behavior in the thermodynamic limit of large domains ($L\gg 1$), which allows some properties to be derived analytically. As in pure active systems [25], the APLG supports phase-separated (PS) stationary states with macroscopic dense (liquid) and dilute (vapor) regions. The densities of active and passive particles in the liquid and vapor phases trace out the binodal curve in the ($\phi_{a},\phi_{p})$ plane, see Fig. 1(a). For $L\gg 1$, this curve can be calculated numerically exactly following the prescription of [36]. This relies on the observation that, for PS states, local concentrations of passive particles and vacancies are proportional
@@
We also compute the spinodal curve as the limit of stability of the homogeneous state (which is $\rho_{\pm}=\phi_{a}/2$, $\rho_{0}=\phi_{p}$). Linear stability analysis of (1) shows that the homogeneous state is unstable when Pe is sufficiently large and $\phi_{p}$ is sufficiently small (see Methods). For the case shown in Fig. 1(a), the dominant eigenvalue of the stability problem is always real, as in the pure active case. Then, the binodal and spinodal are tangent at the critical point (see Fig. 1(a)), which is the standard phenomenology of liquid-vapor phase coexistence for two-component mixtures. Inside the spinodal, the homogeneous state is linearly unstable, and the phase-separated (PS) state is stable; between the spinodal and binodal, the PS state is globally stable while the homogeneous state is metastable.
For larger Pe, this familiar scenario changes qualitatively. In particular, the dominant eigenvalues of the linear stability problem may become complex, leading to dynamical steady states. Such a scenario is illustrated in Fig. 1(b). Crosses on the spinodal mark co-dimension two (Bogdanov–Takens) bifurcations [37] where the dominant eigenvalue becomes complex and the instability of the homogeneous state changes from a pitchfork to a Hopf bifurcation [38]. Crucially, the binodal curve can still be computed for this system: we observe that the spinodal protrudes through the binodal. This effect is intrinsically linked to the existence of nonreciprocal effective interactions between active and passive densities and signals the onset of new physics [9, 22, 15]. (See Appendix D for details of the stability analysis.)
Within the protruding part of the spinodal, the homogeneous state is linearly unstable, and PS states do not exist. It follows that some dynamical steady states must be present in this region. Moreover, in regions where PS states do exist, it may be that (at least) one of the coexisting phases is linearly unstable, which also renders the PS state unstable. The result is that dynamical steady states must exist throughout the blue-shaded region D in Fig. 1(b). Note that this does not rule out the existence of dynamical states elsewhere. Also, while the dominant eigenvalue for the instability is complex for a large part of the spinodal in Fig. 1(b), the resulting steady state may still be (stationary) PS, showing that steady-state properties cannot be deduced directly from linear stability analysis. Dynamical behavior similar to the APLG also occurs in other nonreciprocal systems [9, 22, 15]: we will see below that the APLG provides a specific microscopic model where such generic phenomena can be analyzed quantitatively, including in the thermodynamic limit of large system size.
@@
.3 Illustrative steady states
We present numerical results that illustrate the behavior of the APLG, including direct simulation of the particle model by the Gillespie algorithm [40], and time-stepping the (deterministic) partial differential equation (1) (see Methods). We consider fairly small domains so that the particle model simulations are tractable. [The total number of particles is approximately $(\ell_{y}/\ell_{x})\phi(L/h)^{2}$, such that the simulations of Figs. 2 and 3 involve thousands of particles.]
@@
The binodal construction (Fig. 1) demonstrates the existence of stationary solutions to (1) for large system size $L$. These consist of large $O(L)$ liquid and vapor regions, separated by interfaces of $O(1)$ width. Such phase-separated states are familiar in NRCH, when the nonreciprocity parameter is not too large. In the current setting, the phase separation is caused by the particles’ self-propulsion together with their excluded volume interactions: this is an example of motility-induced phase separation (MIPS) [20, 41, 42, 43], which has been previously characterized in the pure active case ($\phi_{p}=0$) of this model [24].
Fig. 2 shows results of particle-based simulations of the APLG and corresponding numerical solutions of (1) in a domain of size $L=2$. Phase separation occurs in both cases, starting from homogeneous initial states. We find that the dense phase is dominated by active particles, while the dilute phase is mostly passive. In fact, denser phases always have lower concentrations of passive particles because of (4). As usual for MIPS, the magnetization $m$ is large in the interfacial regions but very small within the phases. Note that while the analytic binodal computation shows that stationary PS states exist, these numerical results also show that they are stable (for our choice of parameters).
Figure 4: Traveling profile solution to (1), whose initial condition is the homogeneous state perturbed by a combination of left and right traveling unstable sinusoidal modes and a uniform random perturbation (see Appendix 1, Eqs. (S62,S63) for further details). (a) Density profile at $t=500$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=25$, $\phi_{a}=0.36$, $\phi_{p}=0.3$, $\Delta x=0.05$. Speed of propagation in (6) is $c=-1.8832$.
@@
We now turn to dynamical steady states, which appear in the (blue) region D in Fig. 1(b). We focus on two types: CP solutions retain an overall left-right symmetry with clusters of particles moving in both directions; T solutions break left-right symmetry, leading to a density profile that travels at a fixed speed. Specifically
$\rho_{\sigma}(x,t)=\varrho_{\sigma}(x-ct/L),$ (6)
@@
where $c$ is a constant so the wave velocity is $c/L$; the reason for this $L$-dependence will be discussed below. (There is an analogy of CP and T solutions with standing waves and traveling waves respectively, but note that both CP and T solutions are strongly anharmonic.)
@@
An example CP state is shown in Fig. 3, which again compares direct simulation of the particle model with the numerical solution of (1) whose initial condition is a homogenous state with a uniform random perturbation. This perturbation grows via an instability that involves counterpropagating sinusoidal waves with equal speeds and growth rates. After this transient growth, the resulting time-periodic state consists of two oppositely polarised active clusters that move in opposite directions. As they move, they accumulate passive particles in front of them via a “snowplow effect”. The clusters collide, they pass through each other, and the cycle continues (see Appendix F for additional discussion). In the example of Fig. 3, clusters appear to accelerate during collisions; a simpler example of this effect was observed in [44].
Time-periodic states similar to CP have been observed in other active and nonreciprocal systems [45, 11, 46, 47]: their presence here emphasizes that they are generic. In particular, the APLG still supports the rich dynamical phenomenology of other active-passive mixtures, despite its idealized features. The conection between the APLG and other nonreciprocal systems is discussed further in Appendix B. Despite the analogy with standing waves, we emphasize that these CP clusters experience complex (nonlinear) scattering processes when they meet (see, for example, Fig. S5).
@@
An illustrative T state is shown in Fig. 4, obtained by time-stepping (1) for a larger system ($L=25$). After an initial transient which resembles a CP solution, we find a solution of the form of (6). As in the CP case, this consists of a localized packet of active particles that pushes passive particles in front of it (see Appendix F). We find numerically that such solutions can have a variety of shapes and nontrivial dependence on system size. To simplify this diverse behavior, we again turn to large systems ($L\gg 1$), which enables analytic progress, by analogy with PS states. Hence we reveal new connections between dynamical patterns in this system and equilibrium phase coexistence.
Figure 5: T solutions of the finite-$L$ problem (7) and the large-$L$ problem (8) at $\text{Pe}=7.5$. (Top) Solutions to (8) as a function of $z/L$. (Bottom) Solutions plotted parametrically on the phase diagram of Fig. 1(b), with dashed lines for solutions of (7) and solid lines for solutions of (8). Volume fractions $(\phi_{a},\phi_{p})$ for each column are given in the top, also marked as black triangles in the bottom row. Other parameters: $N=1024$ for numerical solution of (8) and $L=25$, $N=1600$ for (7). Wave speeds for (a,b,c,d) are $c=1.95,1.39,0.41,0.31$ (2 decimal places).
.4 Traveling solutions in large systems